Circular No. 8039

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 Y1 (JUELS-HOLVORCEM) A comet has been found on CCD images (discovery observation below)

taken with a 0.12-m f/5.0 refractor by Charles W. Juels (Fountain Hills, AZ) and Paulo R. Holvorcem (Campinas, Brazil). Holvercem reports that coadding five 45-s and one 90-s exposures taken around Dec. 28.5 UT reveals a coma 1′.8 in diameter. Following posting on the NEO Confirmation Page, several other observers reported on the cometary appearance of the object, including D. T. Durig (Sewanee, TN), who noted that 300-s CCD exposures with a 0.3-m reflector show an inner coma of diameter 30′′ and a wider, diffuse glow of diameter 2′.5.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 28.43892  13^{h}24^{m}01.^{s}30  +16^{o}49′16′′.3  15.1

Additional astrometry and the following very preliminary parabolic orbital elements appear on MPEC 2002-Y50:

T = 2003 Apr. 10.532 TT  ω = 131.^{o}057
Ω = 166.893 }2000.0
q = 0.67088 AU  i = 106.303
2002/03 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 12 12^{h}56.^{m}63 +12^{o}10.^{′}2 2.280 2.194 72.^{o}4 25.^{o}3 16.2
17 13 04.29 +13 17.1 2.134 2.125 76.1 26.7 15.9
22 13 12.47 +14 38.4 1.987 2.055 79.8 28.1 15.6
27 13 21.32 +16 17.1 1.842 1.985 83.4 29.5 15.3
Jan. 1 13 31.02 +18 17.1 1.698 1.915 86.8 30.8 15.0
6 13 41.85 +20 43.0 1.558 1.844 90.1 32.2 14.6
11 13 54.19 +23 40.3 1.424 1.772 93.0 33.7 14.3
16 14 08.66 +27 15.3 1.296 1.700 95.5 35.2 13.9
21 14 26.15 +31 34.3 1.178 1.627 97.2 36.9 13.5

COMETS C/2002 X3 AND C/2002 X4 (SOHO) D. Hammer provides magnitudes from SOHO images (taken with an

orange filter) for the comets reported on IAUC 8032: C/2002 X3, Dec. 4.185 UT, 9.2 ± 0.8; 4.202, 8.5 ± 0.5; 4.213, 8.6 ± 0.6; 4.247, 8.4 ± 0.6. C/2002 X4, Dec. 4.285, 7.3 ± 0.2; 4.309, 7.4 ± 0.2; 4.338, 8.2 ± 0.5; 4.352, 7.9 ± 0.4.

2002 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green